Please solve the question:
One recipe calls for 1/4 teaspoon vanilla. A second recipe requires for 1/2 teaspoon vanilla flour, and a third recipe calls for 1/8 teaspoon vanilla. How much vanilla is required to make all three recipes?

Answer by stanbon(48558)

1/2 + 1/4 + 1/8
In order to add fractions they must all have the same denominator.
So covert 1/2 and 1/4 to have a denominator of 8.
(1/2)(4/4)= 4/8
(1/4)(2/2)= 2/8
Now add 4/8+2/8+1/8=7/8
Cheers,
Stan H.